Calculate next year’s times interest earned ratio, if Cascade sells 2M new shares at $35 per share ($70M) instead of raising new debt:

Next year’s expected EBIT = $115M

Next year’s expected interest expense (without new debt) = $15M

Next year’s expected principal payments (without new debt) = $20M

Common shares outstanding = 25M

Common stock price per share = $40

Company tax rate = 30%

Round intermediate calculations and final answers to the tenth place.
Answer #1

If no new debt are raised, next year interest expenses = $15 Million

Next year expected EBIT = $115 Million

So, next year’s times interest earned ratio = EBIT/(Interest expenses) = 115/15 = 7.67 times